### v3

# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

Abstract
The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many ontology matching (OM) systems. However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on OM tasks at syntactic levels. Our experiments on 8 Ontology Alignment Evaluation Initiative (OAEI) track repositories with 49 distinct alignments indicate: (1) Tokenisation and Normalisation are currently more effective than Stop Words Removal and Stemming/Lemmatisation; and (2) The selection of Lemmatisation and Stemming is task-specific. We recommend standalone Lemmatisation or Stemming with post-hoc corrections. We find that (3) Porter Stemmer and Snowball Stemmer perform better than Lancaster Stemmer; and that (4) Part-of-Speech (POS) Tagging does not help Lemmatisation. To repair less effective Stop Words Removal and Stemming/Lemmatisation used in OM tasks, we propose a novel context-based pipeline repair approach that significantly improves matching correctness and overall matching performance. We also discuss the use of text preprocessing pipeline in the new era of large language models (LLMs).
Index Terms:
ontology matching, text preprocessing
I Introduction
Ontology matching (OM), also known as ontology alignment, is essential to enable interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [1]. A classical OM system usually contains syntactic, lexical, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap [2][3], AgreementMakerLight (AML) [4][5], and FCA-Map [6][7].
Syntactic matching captures “anchor mappings”, providing the foundation for the latter lexical and semantic matching. There are many strategies to extract syntactic information from an ontology entity, including the older approach of Bag-of-Words (e.g. TF-IDF [8]), popular word embedding models (e.g. Word2Vec [9]), and state-of-the-art language models (e.g. BERT [10]). Despite the diversity of the models used, they all apply text preprocessing for cleaning the text data before fitting it into the model. Fig. 1 shows an example of using a classical text preprocessing pipeline to process the ontology entity “cmt:reviewerBiddingStartedBy”. The text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation [11]. Tokenisation is the process of breaking the text into the smallest units (i.e. tokens). We use whitespace to split the tokens in the example. Normalisation is the process of transforming these different tokens into a single canonical form. Stop Words Removal is the process of removing filler words that usually carry little meaning and can be omitted in most cases. Stemming/Lemmatisation is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. This helps to capture the key information from the text and therefore improves efficiency.
Figure 1: Text preprocessing pipeline used in syntactic matching.
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task. Our study tackles the challenge in two ways. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels. In each phase, a text preprocessing method is evaluated for its correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings. Thirdly, we discuss the non-deterministic nature of LLMs for text preprocessing and highlight the importance of retaining the classical text preprocessing pipeline. Specifically, this paper makes the following contributions:
  • (Section IV) We categorise the text preprocessing pipeline used in syntactic OM into two phases. We find a significant improvement using Phase 1 text preprocessing methods. In contrast, Phase 2 text preprocessing methods are currently less effective. We compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation and Lemmatisation + Part-of-Speech (POS) Tagging. We find that inappropriate stop words removal, over-stemming, and over-lemmatisation are common on 8 Ontology Alignment Evaluation Initiative (OAEI) [12] track repositories with 49 distinct alignments.
  • (Section V) We propose a simple and intuitive context-based pipeline repair method. This method is evaluated on the same OM tasks we analysed, showing promising results to improve the correctness of syntactic OM when inserted in the pipeline repair before Phase 2 text preprocessing methods.
  • (Section VI) We discuss the use of text preprocessing pipeline in the era of LLMs. We find LLMs cannot perform syntactic matching without text preprocessing, but they can be used as an alternative to repair the false mappings.
The remainder of the paper is organised as follows: Section II and Section III review the related work and the preliminaries of OM task formulation. Section IV analyses the text preprocessing pipeline used in OM. Section V proposes the context-based pipeline repair approach and experimentally validates its performance. Section VI demonstrates the pros and cons of using LLMs for text processing in OM. Section VII and Section VIII summarise the findings and limitations of this work. Section IX concludes the paper.
II Related Work
Syntactic matching considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology [13]. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge. Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic matching can be traced back to the early stages of OM systems, having been initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker [14] has a normaliser to unify the textual information of the entities. SAMBO [15] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM [16], the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap [17] uses BERT’s inherent WordPiece tokeniser [18] to build the subword of each entity. A more recent approach DeepOnto [19] extends the normalisation to axioms using EL embedding models [20]. The ML extension of LogMap [21] reuses the seed mappings of the traditional system, where each entity is split into its component English word and the mapping is based on their similarity.
However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of text preprocessing on syntactic OM.
III Preliminaries: OM task formulation
Given a source ontology $O_{s}$ and a target ontology $O_{t}$, OM establishes mappings between pairs of entities drawn from each of two ontologies. A correspondence (i.e. one instance of mappings) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence. The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), disjointness ($\perp$), or other more complex relationships. In this paper, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption and disjointness are typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences generated by tools, while a Reference (R) is a set of gold standard correspondences verified by domain experts (i.e. the ground truth alignment).
IV Text Preprocessing Pipeline: Analysis
In this section, we discuss the role of the text preprocessing pipeline in ontology syntactic matching and report our experiments studying the effect of different text preprocessing options over the OAEI repositories.
IV-A Method
Fig. 2 shows the method used to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. Firstly, for both $O_{s}$ and $O_{t}$, we retrieve the entities from the named classes (i.e. owl:Class) and named properties (i.e. object properties owl:ObjectProperty and data type properties owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a numerical identifier), instead, we retrieve the meaningful text from entity labels (i.e. rdfs:label). Then, we apply the text preprocessing pipeline method $f(.)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. Finally, we compare the generated Alignment (A) with Reference (R) to evaluate the performance of the text preprocessing pipeline on OM tasks.
Figure 2: Method used to analyse the text preprocessing pipeline. B-Base Entity without Text Preprocessing, T-Tokenisation, N-Normalisation, R-Stop Words Removal, S/L-Stemming/Lemmatisation.
  1. (1)
    Selected OAEI track repositories: The Ontology Matching Evaluation Toolkit (MELT) [22] is a powerful framework for OM evaluation. We retrieve 17 OAEI normal track repositories stored in the MELT public repository. The repositories are categorised as schema matching or instance matching. 12 of the 17 track repositories are schema-matching and applicable for evaluating OM. Only 8 of these track repositories are selected because the other 4 have noisy data or miss required files. Specifically, the knowledgegraph repository contains both schema and instance matching. The complex repository is primarily focused on detecting complex correspondences, while the multifarm repository is an extension of the conference repository to multilingualism. The laboratory repository lacks reference files at the time of writing. Table I shows the details of the selected track repositories. We consider only the equivalence mappings contained in the reference files. Each track repository may contain more than one alignment corresponding to different pairs of ontologies. For example, the largebio track repository has 6 reference files, pairing the whole and small versions of FMA [23] and NCI [24], FMA and SNOMED [25], and SNOMED and NCI, respectively. The number of references for each track repository is given in the table. There are 49 distinct alignments evaluated in this study.
    Name Domain Number of References anatomy Human and Mouse Anatomy 1 biodiv Biodiversity and Ecology 9 commonkg Common Knowledge Graphs 3 conference Conference 24 food Food Nutritional Composition 1 largebio Biomedical 6 mse Materials Science & Engineering 3 phenotype Disease and Phenotype 2
    TABLE I: Selected OAEI track repositories.
    Compound words are frequently used in ontology naming conventions. For example, compound words “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [26] or the Snake case “art_gallery” used in Wikidata [27]. Fig. 3 shows the proportion of compound words in the selected OAEI track repositories. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 3: The proportion of compound words.
  2. (2)
    Selected subtasks of pipeline methods: There are a variety of subtasks that can be used in a general text preprocessing pipeline, but not all of them are applicable to OM tasks. We select the following subtasks in each text preprocessing method: (1) Tokenisation: includes word tokenisation only. We do not use sentence tokenisation (also known as segmentation) because text retrieved from the entity’s name or label is commonly short. (2) Normalisation: includes lowercasing, HTML tags removal, separator formatting, and punctuation removal. Other subtasks that may potentially change the word semantics, such as removal of special characters and numbers, are excluded. (3) Stop Words Removal: includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [28]. (4) Stemming/Lemmatisation: Stemming methods include Porter Stemmer, Snowball Stemmer, and Lancaster Stemmer. Lemmatisation uses the NLTK Lemmatiser [28] based on WordNet [29], and the word categorisation uses POS Tagging.
  3. (3)
    Selected OM evaluation measures: In information retrieval, there are four primitive measures: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R). Fig. 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
    Figure 4: OM evaluation measures [30].
    Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as:
    $\footnotesize Prec=\frac{|A\cap R|}{|A|}\quad Rec=\frac{|A\cap R|}{|R|}\quad F_{1}=\frac{2}{Prec^{-1}+Rec^{-1}}$ (1)
IV-B Results
  1. (1)
    Comparison of pipeline methods: Fig. 5 summarises the comparison of the text preprocessing methods Tokenization (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The result indicates that the vast majority of correct correspondences are found by T and N. We do not see R and S/L playing a prime role in the OM tasks.
    Figure 5: Comparison of the generic text preprocessing pipeline: Base Entity without Text Preprocessing (B), Tokenisation (T), Normalisation (N), Stop Words Removal (R), Stemming/Lemmatisation (S/L). The methods are always applied sequentially in the pipeline. Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3. The extension of the curves beyond 0% and 100% is an artefact of the seaborn [31] violin plot. (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: the median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
  2. (2)
    Stemming vs. Lemmatisation: Fig. 6 compares Stemming (S) and Lemmatisation (L) in 49 alignments after Tokenisation (T) and Normalisation (N) have been applied. L is commonly better than S.
    Figure 6: Comparison of Stemming (S) and Lemmatisation (L). (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
  3. (3)
    Porter Stemmer vs. Snowball Stemmer vs. Lancaster Stemmer: Fig. 7 compares Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL) in 49 alignments. SP and SS have been found to perform better than SL.
    Figure 7: Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS) and Lancaster Stemmer (SL). (a) Precision: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS. The shape of the distribution is identical.
  4. (4)
    Lemmatisation vs Lemmatisation + POS Tagging: Fig. 8 summarises the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT) in 49 alignments. The result indicates that POS Tagging does not help with Lemmatisation in Precision, Recall, and overall F1 Score in the total of 49 alignments we analysed.
    Figure 8: Comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT). For each of (a) Precision, (b) Recall and (c) F1 Score, the median and the shape of the distribution are identical for L and LT.
IV-C Discussion
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing pipeline methods contain Normalisation and Tokenisation, whereas Stop Words Removal and Stemming/Lemmatisation are assigned to Phase 2. Phase 1 text preprocessing pipeline methods do not change word semantics; instead, they only change syntactic features such as formatting and typography. Conversely, in Phase 2 of the text preprocessing pipeline, words are changed to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
  1. (1)
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method, indicating the benefit of these “rules of thumb” for syntactic OM. If two terms can be matched using a heuristic or intuitive technique, there is no need to leverage the more complex Phase 2 methods. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating for natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e.splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of camel case and snake case.
  2. (2)
    Phase 2 text preprocessing methods: We observe no benefit from Phase 2 text preprocessing methods. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the mapping. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
    1. (a)
      Stop Words Removal: “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and therefore may carry important semantics.
    2. (b)
      Stemming vs. Lemmatisation: Based on our experiments, Lemmatisation is better than Stemming. While lemmatisation tends to avoid generating FPs, it may also miss some implicit TPs. The complexity of finding a missing TP is equal to the size of $(C^{\prime})$. On the other hand, stemming is more aggressive in finding TPs, but the aggression can lead to more FPs as well. The complexity of finding FPs is equal to the smaller size of $(A)$. The workload of discovering FPs after stemming is generally much lighter than detecting missing TPs after lemmatisation, but this may also depend on the accuracy of post-hoc corrections that can be performed.
    3. (c)
      Porter Stemmer vs. Snowball Stemmer vs. Lancaster Stemmer: Porter Stemmer and Snowball Stemmer (also known as Porter 2 Stemmer) have been found to perform better than Lancaster Stemmer, and we cannot see a significant difference between Porter and Snowball. Although The Lancaster Stemmer is more aggressive in detecting more TPs, it does not lead to performance improvement as more FPs are matched.
    4. (d)
      Lemmatisation vs. Lemmatisation + POS Tagging: Lemmatisation + POS Tagging is generally expected to have better results than Lemmatisation alone because tagging can help detect more precise root words. However, we do not observe such a performance improvement when using POS Tagging in our study. The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
V Context-based Pipeline Repair
Experimental results in Section IV demonstrate that only Phase 1 text processing (i.e. Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (i.e. Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour. Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets.
V-A Causal analysis
  1. (1)
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only helps detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on Precision, Recall, and overall F1 Score.
    $\footnotesize Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$ (2)
    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (3)
    $\footnotesize F_{1}\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$ (4)
    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties [32] [33].
  2. (2)
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meaning, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, Recall is always increasing, but Precision and overall F1 Score are less reliable.
    $\footnotesize Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$ (5)
    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (6)
    $\footnotesize F_{1}\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$ (7)
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the threshold to increase Precision and F1 Score is:
    $\scriptsize\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$ (8)
    In our experiments, we actually observe a reduction in Precision and overall F1 Score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post-hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
V-B Repair Method
Phase 2 text preprocessing pipeline methods (Stop Words Removal and Stemming/Lemmatisation) have been shown to be less effective in OM tasks, caused mainly by FPs. We propose a pipeline repair approach that aims to differentiate FPs and therefore improve Precision and F1 Score. As illustrated in Fig. 9: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while FPs decrease. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but FPs also increase. (3) Our context-based pipeline repair approach collapses Alignment (A). The number of FPs significantly decreases, with a slight decrease in TPs.
Figure 9: Mechanism of the context-based pipeline repair.
One critical step in our approach is to retrieve a reserved word set that may cause FPs after the text preprocessing, and these words will be excluded before the text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (we call our approach context-based here because pairs of words may have the same or different meanings in different contexts). Based on these two assumptions, we propose a simple and intuitive Algorithm 1 to retrieve the reserved word set for context-based pipeline repair. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in each of them. For a specific text preprocessing method $f(.)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we also remove the words where $f(w)=w$ from the final set because either skipping or keeping these words in the reserved word set would not change the mapping results.
Algorithm 1 Finding the reserved words
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
0:  $Reserved\_Word\_Set$
  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
  /* Phase 1.1: Find duplicates in $O_{s}$/
  for $Entity\ e_{i},e_{j}\in O_{s}$ do
     if $f(e_{i})=f(e_{j})$ then
        $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
     end if
  end for
  /* Phase 1.2: Find duplicates in $O_{t}$/
  /* Same procedure applies …*/
  /* Phase 2: Find duplicates in $Reserved\_Word\_Set$/
  for $Word\ w\in Reserved\_Word\_Set$ do
     if $f(w)=w$ then
        $Reserved\_Word\_Set\rightarrow w$
     end if
  end for
  return  $Reserved\_Word\_Set$
An example of generating and using a simple reserved word set is illustrated below. (1) Two object properties was_a_member_of and has_members from a single ontology have the same result “member” via the traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are duplicates in the reserved set. We can observe that the word “member” is a duplicate because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are preprocessed as “has a steer committe” and “was a steer committe of”, respectively.
V-C Evaluation
We apply our context-based pipeline repair approach to the same OAEI track repositories and alignments as above. Fig. 10 compares with and without context-based pipeline repair in 8 track repositories with 49 distinct alignments. The text preprocessing pipeline methods implemented include all the Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT). We can see that our context-based repair approach can significantly improve the Precision. As a trade-off, it may cause a slight decrease in Recall, but the overall F1 Score is still increasing in the majority of the alignments. For example, this repair approach applied to the MaterialInformation-EMMO alignment using Lancaster Stemmer improved Precision and overall F1 score by 8.95% and 2.92%, with only a decrease of 3.17% in Recall.
Figure 10: Context-based pipeline repair use in Phase 2 text preprocessing methods (the total number of each category can appear to be less than 49 when data points overlap).
VI Text Preprocessing in the era of LLMs
With the recent advances in LLMs, we ask whether we still need the classic text processing pipeline. In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the benefits of LLMs in a repair function inside the OM text preprocessing pipeline.
VI-A LLMs Used in OM Text Preprocessing
We observe that some LLMs can perform syntactic matching without preprocessing, but their robustness remains questionable. We conduct a study on the use of LLMs for syntactic matching with and without text preprocessing. The experiment settings are described as follows:
  1. (1)
    We select the same example described in Section IV-A. The compound word “art gallery” can have different naming conventions in different ontologies. For example, the Pascal case “ArtGallery” used in YAGO and the Snake case “art_gallery” used in Wikidata.
  2. (2)
    We choose 10 LLMs from 4 different families. These include two OpenAI models (gpt-3.5-turbo and gpt-4-turbo) [34], three Anthropic Claude models (claude-3-opus, claude-3-sonnet, and claude-3-haiku) [35], three Mistral AI models (mistral-large, mistral-medium, and mistral-small) [36], and two open-source Meta Llama models (llama-3-70b and llama-3-8b) [37]. We use LangChain [38] to build the chats with OpenAI, Mistral AI, and Anthropic Claude models, and Ollama [39] to access Meta Llama models. The versions of LLMs used in the experiment are listed in Table II. All the model temperatures are set to 0 to minimise the random results generated by LLMs.
    Series LLMs Versions GPT gpt-4-turbo gpt-4-turbo-2024-04-09 gpt-3.5-turbo gpt-3.5-turbo-0125 Claude claude-3-opus claude-3-opus-20240229 claude-3-sonnet claude-3-sonnet-20240229 claude-3-haiku claude-3-haiku-20240307 Mistral mistral-large mistral-large-2402 mistral-medium mistral-medium-2312 mistral-small mistral-small-2402 Llama llama-3-70b date accessed: 2024-06-01 llama-3-8b date accessed: 2024-06-01
    TABLE II: Versions of LLMs used in the experiment.
  3. (3)
    We generate the prompt template as “Is ArtGallery <KEYWORD> to/with art_galley?” (without text preprocessing) and “Is art galley <KEYWORD> to/with art galley?” (with text preprocessing). We use only T and N to preprocess the text because R and S/L are shown to be less effective in the main content of the paper. Furthermore, R and S/L needs are not evident in the original naming of the ontology entities. The <KEYWORD> is a placeholder for a collection of words that can be used to describe an equivalence relationship. In this study, we experiment with three common words namely “identical”, “interchangeable”, and “equivalent”.
  4. (4)
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt). We use parentheses to indicate that the different responses generated by the prompts have a self-reflection phase.
Table III shows the results of using LLMs for syntactic matching without text preprocessing. We expect the LLMs to generate “Yes” answers for this TP sample. However, the majority of the LLMs output the “No” answers (marked with the red colour in the table) without text preprocessing.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo No No No gpt-3.5-turbo No No No claude-3-opus No Yes Yes claude-3-sonnet No No No claude-3-haiku No No No mistral-large No No Yes mistral-medium No No No mistral-small No Yes Yes (No) llama-3-70b No No No llama-3-8b Yes Yes Yes
TABLE III: Using LLMs for syntactic matching without text preprocessing. Prompt: “Is ArtGallery <KEYWORD> to/with art_gallery? Answer yes or no. (Write a short explanation.)”
Table IV shows the results of using LLMs for syntactic matching with text preprocessing. We can see that the error rate can be significantly reduced, but some LLMs still continue to produce incorrect results. The result also shows that LLMs with a larger number of parameters (e.g. llama-3-70b) and prompts with a self-reflection phase (labelled in parentheses) offer “too much of a good thing”. They do not help with syntactic matching and may even produce more errors in OM tasks. Although this preliminary study does not cover all possible triggers used in prompt engineering, it is adequate to highlight the importance of proper preprocessing before fitting the text into the LLMs. We believe that the classical programming-based text preprocessing pipeline is more stable and reliable than using LLM’s implicit text preprocessing or prompt-based text preprocessing for OM tasks.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo Yes Yes Yes gpt-3.5-turbo Yes Yes Yes claude-3-opus Yes Yes Yes claude-3-sonnet Yes (No) No Yes (No) claude-3-haiku No No Yes (No) mistral-large Yes Yes Yes mistral-medium No Yes Yes mistral-small Yes Yes Yes llama-3-70b Yes Yes (No) Yes (No) llama-3-8b Yes Yes Yes
TABLE IV: Using LLMs for syntactic matching with text preprocessing. Prompt: “Is art gallery <KEYWORD> to/with art gallery? Answer yes or no. (Write a short explanation.)”
VI-B LLMs Used in OM Text Preprocessing Pipeline Repair
Using LLMs for OM text preprocessing is not without merit. With strong background knowledge, they can be used to facilitate several subtasks in OM. For example, LLMs can be used as a repair tool for the text preprocessing pipeline, as an alternative to the approach we proposed in Section V-B. The slight difference is that the context-based pipeline repair approach is an ad hoc repair prior to syntactic matching, while the LLM-based approach is a post hoc repair after syntactic matching. We conduct a preliminary study on LLMs used in OM text preprocessing pipeline repair. The experiment settings are described as follows:
  1. (1)
    The experiment is set up on the largebio track repository, a biomedical track repository requires higher Precision with low rates of FPs. The Lancaster Stemmer always shows lower Precision across 6 pairs of alignments, meaning that the results contain a significant number of FPs.
  2. (2)
    We choose the same 10 LLMs used in the previous experiment in Appendix VI-A.
  3. (3)
    We use the prompt “Is X equivalent to Y?” and preprocess the entity names X and Y with N and T before fitting them into the LLMs. The previous experiment in Appendix VI-A shows these two settings are more suitable for OM tasks.
  4. (4)
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt).
Figure 11: The discovery rate of FPs using LLMs on the largebio track repository.
Fig. 11 shows the discovery rate of FPs using 10 different LLMs. We expect the LLMs to generate “No” answers for these FP samples. The GPT model gpt-4-tubo, the Claude model claude-3-haiku, the Mistral model mistral-medium, and the Llama model llama-3-70b achieve better performance in detecting FPs. For the GPT and Llama series, LLMs with a larger number of parameters perform better than those with a relatively smaller number of parameters. For the Claude and Mistral series, LLMs with larger parameters do not improve performance. In the test cases we analysed, they are even less effective than those with a relatively smaller number of parameters. Unlike detecting TPs, we can see a self-reflection phase improves the overall performance of detecting FPs.
VI-C Implications
LLMs were originally developed for the purpose of question-answering (QA). OM is not only a QA task, but also a knowledge-intensive task that relates to a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities of using LLMs for OM tasks. From our experience, we highlight two pathways to customise LLM for a new domain task. One pathway is function calling (FC), which refers to teaching LLMs using external tools. In LLM-based text preprocessing, we could package the programming-based text preprocessing pipeline as a tool for LLMs to use. The other pathway is fine-tuning (FT), where LLMs are customised with local data. In LLM-based text preprocessing repair, we could use human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs, so that they can better understand the context of the task.
VII Discussion
Ontologies provide formalised conceptualisations of knowledge graphs (KGs). Incorporating ontologies for KG-related NLP tasks can enhance the capability to handle complex concepts and relations, thereby improving the quality of KGs built from text. Ontologies also assist interoperability of text generated from KGs. However, aligning and integrating heterogeneous ontologies remains a challenge when using them for NLP tasks. OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
OM is a complex process that combines syntactic, lexical, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This paper concentrates on text preprocessing, a common practice in syntactic matching. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experience, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
  1. (1)
    Our study shows (i) whether to use, or not use, and (ii) how to use these text preprocessing methods. Such experimental results will benefit decision making when selecting appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
  3. (3)
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with syntactic, lexical, and semantic matching. LLMs may not outperform traditional methods in performing these sub-matching tasks. The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer readers to several recent works in [40, 41, 42, 43, 44].
This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the overall performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
VIII Limitations
OAEI track repositories are comprehensive but do not cover all real-world scenarios. Furthermore, the reference files contained in the OAEI track repositories are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between named classes, object properties, and data type properties. Disjointness mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI track repositories contain subsumption mappings, and they are not enough to robustly demonstrate text preprocessing used in subsumption mappings. This work has not been entered into the official OAEI campaign because of its intense focus on only the preprocessing step.
IX Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI track repositories with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. Recall) and correctness (i.e. Precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. Recall), but matching correctness (i.e. Precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. Precision) and overall matching performance (i.e. F1 Score). (3) We also argue the traditional text preprocessing pipeline is not obsolete in the era of LLMs, but an integral part for LLMs used in the OM tasks.
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline for OM tasks. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Supplemental Material Statement
The source code for this study is available from https://github.com/qzc438-research/ontology-nlp. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment
The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.