### v5

# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching? [Experiment, Analysis & Benchmark]

Abstract.
The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). To repair the unwanted false mappings caused by Phase 2 text preprocessing, we propose a novel context-based pipeline repair approach that employs a post hoc check to find common words that cause false mappings. These words are stored in a reserved word set and applied in text preprocessing. The experimental results show that our approach improves the matching correctness and the overall matching performance. We then consider the broader integration of the classic text preprocessing pipeline with modern large language models (LLMs) for OM. We recommend that (1) the text preprocessing pipeline be injected via function calling into LLMs to avoid the tendency towards unstable true mappings produced by LLM prompting; or (2) LLMs be used to repair non-existent and counter-intuitive false mappings generated by the text preprocessing pipeline.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-nlp.
1. Introduction
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classical OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021). While syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching, there are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before using the model. The pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM. T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas (is)Reviewing
Table 1. Examples of using the text preprocessing pipeline in syntactic OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word form.
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies question its role in optimising the matching performance. Our study fills this gap. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM. This method shows promising results for repairing less effective text preprocessing methods. Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classic text preprocessing pipeline with LLM-based OM. Specifically, this paper makes the following contributions:
  • (Section 4) We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.d.]) datasets with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find inappropriate stop words removal, over-stemming, and over-lemmatisation. We also compare the performance of several text preprocessing methods: (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
  • (Section 5) We propose a simple and intuitive context-based pipeline repair method using a reserved word set to capture certain words that may lead to false mappings. This method is evaluated on the same OAEI datasets we analysed, showing promising results in improving the matching correctness and the overall matching performance of syntactic OM.
  • (Section 6) We discuss the integration of the classic text preprocessing pipeline with modern LLMs. We recommend using text preprocessing via function calling to overcome the non-determinism of purely prompt-based approaches to finding true mappings, while LLMs show great potential to repair false mappings generated by the text preprocessing pipeline.
The remainder of the paper is organised as follows. Section 2 and 3 review the related work and the preliminaries of OM task formulation. Section 4 analyses the text preprocessing pipeline used in syntactic OM. Section 5 proposes the context-based pipeline repair approach and experimentally validates its performance. Section 6 demonstrates the integration of the classic text preprocessing pipeline with modern LLMs in syntactic OM. Section 7 and 8 discuss the findings and limitations. Section 9 concludes.
2. Related Work
Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
3. Preliminaries: OM Task Formulation
Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships. In this study, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e. the ground truth alignment).
4. Text Preprocessing Pipeline Analysis
In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options on the OAEI datasets.
4.1. Experiment Setup
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
  1. (1)
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study.
    Table 2. Selected OAEI tracks and the number of alignments.
    Name Domain Number of Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
    Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 2 shows the proportion of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 2. Proportion of compound words.
  2. (2)
    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM. We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words. Experiments on several subword tokenisation methods (e.g. Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e. converting the tokens to a sentence) are generally the same as full word tokenisation. We do not use sentence tokenisation (a.k.a. segmentation) because text retrieved from the entity’s name or label is commonly short. (b) Normalisation (N): includes lowercasing, HTML tag removal, separator formatting, and punctuation removal. Other methods that may potentially change the word semantics are excluded (e.g. removal of special characters and numbers). (c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008). (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster. Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
  3. (3)
    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts. Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\{(e_{1},e_{2},\equiv)|e_{1}\in O_{s},e_{2}\in O_{t}\}$. Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as follows.
    Figure 3. OM Evaluation (Euzenat, 2007).
    $Precision(Prec)=\frac{|A\cap R|}{|A|}$
    $Recall(Rec)=\frac{|A\cap R|}{|R|}$
    $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
4.2. Results
4.2.1. Comparison of Text Preprocessing Pipeline
Figure 4 compares the stages of the text preprocessing pipeline. Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The methods are always applied sequentially in the pipeline. For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM. For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role. Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3. The extension of the curves beyond 0% and 100% is an artefact of the Seaborn (Waskom, 2021) package. In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: The median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
Figure 4. Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The symbol $\varnothing$ represents the baseline without text preprocessing.
4.2.2. Stemming vs. Lemmatisation
Figure 5 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied. Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded). In the violin plots: (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
Figure 5. Stemming (S) vs. Lemmatisation (L).
4.2.3. Different Stemming Methods
Figure 6 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
Figure 6. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
4.2.4. Lemmatisation with/without POS Tagging
Figure 7 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
Figure 7. Lemmatisation (L) without/with POS Tagging (LT).
4.3. Implications
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation. Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography. In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
  1. (1)
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
  2. (2)
    Phase 2 text preprocessing methods: We observe no benefit in using Phase 2 text preprocessing methods in syntactic OM. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
    1. (a)
      Stop Words Removal (R): Logical expressions “and”, “or”, and “not” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics. Some single stop word characters (e.g. “d”, “i”, “m”, “o”, “s”, “t”, and “y”) should be removed from the stop word set because contractions (e.g. I’d, I’m, o’clock, it’s, isn’t, y’know) are unlikely to appear in ontology entities, and yet those single characters are likely to carry important meaning if they do occur (e.g. “Chromosome_Y”).
    2. (b)
      Stemming (S) vs. Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found. Based on our experiments with no post hoc correction applied, L is better than S.
    3. (c)
      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball. Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
    4. (d)
      Lemmatisation (L) vs. Lemmatisation + POS Tagging (LT): LT is generally expected to have better results than L alone because tagging can help detect more precise root words. However, we do not observe such a performance improvement when using POS tagging in our study. The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
  3. (3)
    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
5. Context-Based Pipeline Repair
The experimental results in Section 4 demonstrate that only Phase 1 text processing (Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour and introduce an intuitive context-based pipeline repair approach.
5.1. Causal Analysis
  1. (1)
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
    (1) $Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$
    (2) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (3) $F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$
  2. (2)
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
    (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
    (5) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
    (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
    In our experiments, we actually observe a reduction in precision and overall F1 score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post hoc correction phase in which removing FPs is considered to be an easier human task than adding missing TPs. This idea inspires our novel pipeline repair method.
5.2. Repair Method
While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Figure 8: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our context-based pipeline repair aims to collapse Alignment (A). The number of FPs decreases significantly, with a slight decrease in TPs.
Figure 8. Motivation of our context-based pipeline repair.
One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies. We call our approach context-based here because pairs of words may have the same or different meanings in different contexts. Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
Algorithm 1 Finding Reserved Word Set
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
0:  $Reserved\_Word\_Set$
  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
  /* Phase 1.1: Find duplicates in $O_{s}$/
  for $Entity\ e_{i},e_{j}\in O_{s}$ do
     if $f(e_{i})=f(e_{j})$ then
        /* Pick up every different word between $e_{i}$ and $e_{j}$*/
        $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
     end if
  end for
  /* Phase 1.2: Find duplicates in $O_{t}$/
  /* Same procedure applies …*/
  /* Phase 2: Find immutable words in $Reserved\_Word\_Set$/
  for $Word\ w\in Reserved\_Word\_Set$ do
     if $f(w)=w$ then
        $Reserved\_Word\_Set\rightarrow w$
     end if
  end for
  return  $Reserved\_Word\_Set$
An example of generating and using a simple reserved word set is illustrated below. (1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set. We can observe that the word “member” is immutable because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
5.3. Evaluation
We apply our context-based pipeline repair approach to the same OAEI datasets as above. Figure 9 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments. The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
Figure 9. Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
  1. (1)
    Precision: Most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness. Recall: Most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness. F1 Score: Most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
  2. (2)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
6. Text Preprocessing Pipeline & LLMs
In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classic text preprocessing pipeline with modern LLMs.
6.1. Motivation
In modern LLMs, text preprocessing can be performed implicitly within the models. For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats. However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classic text processing pipeline. On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline. For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classic text preprocessing method stemming cannot.
6.2. Experiment Setup
  1. (1)
    We choose 10 LLMs from 4 different families. These include two OpenAI GPT models (OpenAI, [n.d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.d.]), and ChatGLM 4 (Zeng et al., 2024). We use LangChain (LangChain, Inc., [n.d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.d.]) to access open-source models. Table 3 lists the details of selected LLMs. All model temperatures are set to 0 to minimise the random results generated by LLMs.
    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
    Family Model Version Type GPT gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 $\blacktriangle$ Claude claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 $\blacktriangle$ Llama llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$ llama-3.1-8b Ollama Model ID: 46e0c10c039e $\bullet$ Qwen qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$ qwen-2.5-7b Ollama Model ID: 845dbda0ea48 $\bullet$ Gemma gemma-2-9b Ollama Model ID: ff02c3702f32 $\bullet$ GLM glm-4-9b Ollama Model ID: 5b699761eca5 $\bullet$
  2. (2)
    We select 4 different LLM prompt strategies. (a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2? Answer yes or no.” (b) Few-Shot: Adding an example in the prompt, such as “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation. We add a sentence “Write a short explanation.” to invoke this function. (d) Few-Shot + Self-Reflection: A combination of (b) and (c). Table 4 lists the details of the selected prompt strategies and their templates.
    Table 4. Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
    No. Prompt Strategy Prompt Template PT1 Basic Is Entity1 equivalent to Entity2? Answer yes or no. PT2 Few-Shot Example: Hair_root is equivalent to Hair_Root. Is Entity1 equivalent to Entity2? Answer yes or no. PT3 Self-Reflection Is Entity1 equivalent to Entity2? Answer yes or no. Write a short explanation. PT4 Few-Shot Example: Hair_root is equivalent to Hair_Root. + Is Entity1 equivalent to Entity2? Answer yes or no. Self-Reflection Write a short explanation.
  3. (3)
    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
  4. (4)
    We report our results over a single run. There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
6.3. Detecting True Mappings
LLMs are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, we employ the pipeline via function calling (OpenAI, 2023) (a.k.a. tool calling) in LLMs. This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method. Two entities are considered to be matched if they have the same results after the text preprocessing.
Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly caused by “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
6.4. Detecting False Mappings
LLMs are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs are non-trivial mappings that cannot be detected by the classic text preprocessing method.
Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
6.5. Implications
LLMs were originally developed and trained for question answering (QA). OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM. We highlight two pathways for this integration. One pathway is to use function calling to invoke the classic text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use. This provides stable results for text preprocessing in syntactic OM. The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction to reduce the number of FPs. This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and to feed the validated data back into LLMs so that they can better understand the context of the task (i.e. fine-tuning LLMs for syntactic OM).
7. Discussion
OM aims to bridge the gap by capturing similar concepts that occur in different ontologies. OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
  1. (1)
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical matching and semantic matching.
  3. (3)
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with different matching techniques. LLMs may not outperform traditional methods in performing these sub-matching tasks. We suggest “rules of thumb” for syntactic OM. If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose. The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
  4. (4)
    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
8. Limitations
The references contained in the OAEI tracks are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings. This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
9. Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score). (3) We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs. It is an integral part of the LLMs used in OM tasks. Integrating the classic text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches. Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Acknowledgements.
The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets. The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track. The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project. According to the OAEI data policy (retrieved January 1, 2025), “OAEI results and datasets, are publicly available, but subject to a use policy similar to the one defined by NIST for TREC. These rules apply to anyone using these data.” Please find more details from the official website: https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.