### v7

# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

Abstract.
The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing. One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings. These words are stored in a reserved word set and applied before the text preprocessing. By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach. This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing. It also overcomes the tendency towards unstable true mappings by injecting the classic text preprocessing pipeline via function calling. The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
Ontology Matching, Text Preprocessing
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
1. Introduction
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classic OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
Syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching. There are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
The classic text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM. In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
No. Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
Table 1. Examples of using the classic text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance. Our study fills this gap. We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches. These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline. Specifically, this paper makes the following contributions:
  • We compare the performance of the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n. d.]) with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
  • We propose two approaches to repair the less effective text preprocessing pipeline in syntactic OM: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair. Experimental results show that these two methods can improve the matching correctness and the overall matching performance of syntactic OM.
The remainder of the paper is organised as follows. Section 2 and 3 review the related work and the preliminaries of OM and syntactic matching. Section 4 analyses the text preprocessing pipeline used in syntactic OM. Section 5 proposes text preprocessing pipeline repair approaches and experimentally validates their performance. Section 6 highlights the contributions of this work and concludes the paper.
2. Related Work
Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
Early explorations in (He et al., 2023; Norouzi et al., 2023) show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance. OLaLa (Hertling and Paulheim, 2023) proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution. Works on both Agent-OM (Qiang et al., 2023) and LLM4OM (Giglou et al., 2024) point out that LLMs have limited context length and inherited hallucinations. While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling (OpenAI, 2023) for OM to improve matching performance. The authors in (Taboada et al., 2025) state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection. Recent works in (Amini et al., 2024; Zamazal, 2024; Sousa et al., 2025) also utilise LLMs and LLM embeddings for complex OM. However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for syntactic matching. For syntactic OM, the challenges of using LLMs are twofold. On one hand, LLM hallucinations may produce incorrect “seed mappings”. Such mappings are toxic for the subsequent lexical and semantic matching. On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
3. Preliminaries
An ontology provides a formalised conceptualisation of domain knowledge (Gruber, 1993). The basic component of an ontology is the entity. Each entity is assigned a unique identifier (i.e. IRI) and represents a unique concept of the domain. Entities can be a concept (i.e. class), relation (i.e. property), or instance (i.e. individual) (Noy et al., 2001). Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap. It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names (OBO Foundry, [n. d.]).
Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships (Euzenat and Shvaiko, 2013). OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. We focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
4. Text Preprocessing Pipeline Analysis
In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options on the OAEI datasets.
4.1. Experiment Setup
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. While OM methods will often follow text preprocessing with a more sophisticated strategy to determine equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone. This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
[…]
Chart of experiment setup to analyse the text preprocessing pipeline.
  1. (1)
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study. Figure 2 shows the number of entities in each alignment.
    Table 2. Selected OAEI tracks and the number of alignments.
    Name Domain Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
    Figure 2. Number of entities in each alignment.
[…]
    Chart of number of entities in each alignment.
    Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 3(b) shows the frequency distribution of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 3. Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
[…]
    Chart of frequency distribution of compound words.
  2. (2)
    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM. We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words. Experiments on several subword tokenisation methods (e.g. Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e. converting the tokens to a sentence) are generally the same as full word tokenisation. We do not use sentence tokenisation (a.k.a. segmentation) because text retrieved from the entity’s name or label is commonly short. (b) Normalisation (N): includes lowercasing, HTML tag removal, separator formatting, and punctuation removal. Other methods that may potentially change the word semantics are excluded (e.g. removal of special characters and numbers). (c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008). (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster. Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
  3. (3)
    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts. Figure 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences. Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as follows.
    Figure 4. OM evaluation (Euzenat, 2007).
[…]
    Chart of OM evaluation (Euzenat, 2007).
    $Precision(Prec)=\frac{|A\cap R|}{|A|}$
    $Recall(Rec)=\frac{|A\cap R|}{|R|}$
    $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
4.2. Results
Figure 5 compares each step of the text preprocessing pipeline. These steps are always applied sequentially in the pipeline. Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM. For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role. In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: The median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
Figure 5. Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn (Waskom, 2021) package).
[…]
Comparison of text preprocessing pipeline.
  1. (1)
    Figure 6 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied. Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded). In the violin plots: (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
    Figure 6. Stemming (S) vs. Lemmatisation (L).
[…]
    Stemming (S) vs. Lemmatisation (L).
  2. (2)
    Figure 7 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
    Figure 7. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
[…]
    Comparison of different stemmers.
  3. (3)
    Figure 8 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
    Figure 8. Lemmatisation (L) without/with POS Tagging (LT).
[…]
    Lemmatisation (L) without/with POS Tagging (LT).
4.3. Implications
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation. Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography. In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
  1. (1)
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
  2. (2)
    Phase 2 text preprocessing methods: We observe no benefit in using Phase 2 text preprocessing methods in syntactic OM. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
    1. (a)
      Stop Words Removal (R): Logical expressions “and”, “or”, and “not” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics. Some single stop word characters (e.g. “d”, “i”, “m”, “o”, “s”, “t”, and “y”) should be removed from the stop word set because contractions (e.g. I’d, I’m, o’clock, it’s, isn’t, y’know) are unlikely to appear in ontology entities, and yet those single characters are likely to carry important meaning if they do occur (e.g. “Chromosome_Y”).
    2. (b)
      Stemming (S) vs. Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found. Based on our experiments with no post hoc correction applied, L is better than S.
    3. (c)
      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball. Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
    4. (d)
      Lemmatisation (L) vs. Lemmatisation + POS Tagging (LT): LT is generally expected to have better results than L alone because tagging can help detect more precise root words. However, we do not observe such a performance improvement when using POS tagging in our study. The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
  3. (3)
    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
5. Text Preprocessing Pipeline Repair
The experimental results in Section 4 demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour and introduce our novel pipeline repair approaches.
5.1. Causal Analysis
  1. (1)
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
    (1) $Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$
    (2) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (3) $F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$
  2. (2)
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
    (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
    (5) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
    (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
    In our experiments, we actually observe a reduction in precision and overall F1 score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post hoc correction phase in which removing FPs is considered to be an easier human task than adding missing TPs. This idea inspires our novel pipeline repair method.
5.2. Motivation
While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Figure 9: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our text preprocessing pipeline repair aims to collapse Alignment (A). The number of FPs decreases significantly, with a slight decrease in TPs.
Figure 9. Motivation of our context-based pipeline repair.
[…]
Motivation of our context-based pipeline repair.
5.3. Ad Hoc Logic-Based Repair
The ad hoc logic-based repair applies before the text preprocessing. One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section 3). This approach is context-based because pairs of words may have the same or different meanings in different contexts. Based on these two features, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
Algorithm 1 Finding Reserved Word Set
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(\cdot)$
0:  Reserved_Word_Set $S$
  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
  /* Find duplicates in $O_{s}$/
  for $Entity\ e_{i},e_{j}\in O_{s}$ do
     if $f(e_{i})=f(e_{j})$ then
        /* Pick up every different word between $e_{i}$ and $e_{j}$*/
        $diff(e_{i},e_{j})=(e_{i}\setminus e_{j})\cup(e_{j}\setminus e_{i})$
        $S=S+diff(e_{i},e_{j})$
     end if
  end for
  /* Find duplicates in $O_{t}$/
  /* Same procedure applies …*/
  /* Phase 2: Find immutable words in $S$/
  for $Word\ w\in S$ do
     if $f(w)=w$ then
        $S=S-w$
     end if
  end for
  return  Reserved_Word_Set $S$
An example of generating and using a simple reserved word set is illustrated below. (1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set. We can observe that the word “member” is immutable because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
5.4. Post Hoc LLM-Based Repair
With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing. However, there is a wide range of LLMs with different prompting strategies that can be used in LLM-based repair. We aim to find a generic pipeline repair approach and the best practice to integrate LLMs in the repair phase. The experiment settings are as follows.
  1. (1)
    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
  2. (2)
    We choose 10 LLMs from 4 different families. These include two OpenAI GPT models (OpenAI, [n. d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n. d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n. d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n. d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n. d.]), and ChatGLM 4 (Zeng et al., 2024). We use LangChain (LangChain, Inc., [n. d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n. d.]) to access open-source models. Table 3 lists the details of selected LLMs. All model temperatures are set to 0 to minimise the random results generated by LLMs.
    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
    Family Model Version GPT $\blacktriangle$ gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 Claude $\blacktriangle$ claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 Llama $\bullet$  llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$  llama-3.1-8b Ollama Model ID: 46e0c10c039e Qwen $\bullet$  qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$  qwen-2.5-7b Ollama Model ID: 845dbda0ea48 Gemma $\bullet$  gemma-2-9b Ollama Model ID: ff02c3702f32 GLM $\bullet$  glm-4-9b Ollama Model ID: 5b699761eca5
  3. (3)
    We select 4 different LLM prompt strategies. (a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2? Answer yes or no.” (b) Few-Shot: Adding an example in the prompt. We use the example “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation. We add a sentence “Write a short explanation.” to invoke this function. (d) Few-Shot + Self-Reflection: A combination of (b) and (c). Table 4 lists the details of the selected prompt strategies and their templates.
    Table 4. Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
    No. Prompt Strategy Prompt Template PT1 Basic Is Entity1 equivalent to Entity2? Answer yes or no. PT2 Few-Shot Example: Hair_root is equivalent to Hair_Root. Is Entity1 equivalent to Entity2? Answer yes or no. PT3 Self-Reflection Is Entity1 equivalent to Entity2? Answer yes or no. Write a short explanation. PT4 Few-Shot Example: Hair_root is equivalent to Hair_Root. + Is Entity1 equivalent to Entity2? Answer yes or no. Self-Reflection Write a short explanation.
  4. (4)
    We compare LLMs with the classic text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
  5. (5)
    We report our results over a single run. There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
Figure 10(l) compares LLMs with the classic text preprocessing pipeline in detecting TPs. LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
[…]
Discovery rate of TPs using LLMs on the Largebio Track.
Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
[…]
Discovery rate of FPs using LLMs on the Largebio Track.
Figure 11(l) compares LLMs with the classic text preprocessing pipeline in detecting FPs. LLMs and the text preprocessing pipeline are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classic text preprocessing method. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with
  1. (1)
    Employ the classic text preprocessing pipeline via function calling (a.k.a. tool calling) to detect TPs. This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method. This provides stable results for text preprocessing when finding FPs. LLM prompting is not suitable for this task, even with complex prompting strategies.
  2. (2)
    Use LLM prompting to detect FPs. The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
5.5. Evaluation
We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL), and evaluate their performance on the three large datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
Figure 12(c) compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR). We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases. Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score. For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching. These mappings are meant to be precise and accurate. Both of our pipeline repair approaches have shown high precision and are the best tool suites.
(a) WHOLE-FMA-NCI
(b) WHOLE-FMA-SNOMED
(c) WHOLE-SNOMED-NCI
Figure 12. Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR). The symbol $\varnothing$ represents the baseline without text preprocessing repair.
[…]
Evaluation of pipeline repair approaches.
Figure 13 evaluates the scalability of ad hoc logic-based repair (AR). The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities. For all the alignments we tested, the value of reserved words @100 is less than 10. With the number of entities increasing (i.e. the ontology size increases), the value of reserved words @100 does not increase but decreases. This feature of our approach is scalable in large-scale syntactic OM.
Figure 13. Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
[…]
Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
Figure 14(c) compares post hoc LLM-based repair (PR) with two alternative repair approaches. The results show that PR achieves almost the same precision with text preprocessing only repair (TOR) and LLM only repair (LOR), but its recall and F1 score are better than the other two alternatives. For syntactic OM, this means that our approach could detect more ”seed mappings”.
(a) WHOLE-FMA-NCI
(b) WHOLE-FMA-SNOMED
(c) WHOLE-SNOMED-NCI
Figure 14. Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches: text preprocessing only repair (TOR) and LLM only repair (LOR).
[…]
Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
6. Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings. Both approaches outperform the traditional text preprocessing pipeline without repair, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score).
While the classic text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
  1. (1)
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
    Our pipeline repair approaches can be used to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
  3. (3)
    We showcase the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM. We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic. The best practice is to integrate the classic text preprocessing pipelines with the LLMs. Using function calling to invoke the classic text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
  4. (4)
    Our study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Acknowledgements.
The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets. The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track. The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
Limitations
The references contained in the OAEI tracks are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings. This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
Artifacts
The source code is available from Github at https://github.com/qzc438-research/ontology-nlp. The OAEI datasets are available from the OAEI Matching EvaLuation Toolkit (MELT) at https://dwslab.github.io/melt/track-repository (retrieved January 1, 2025). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Appendix A Missing References in the LargeBio Track
As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may miss some reference alignments. Here we list the missing reference alignments found by the text preprocessing pipeline in the folder “csvs”. These mappings have outstanding possibilities of being correct, and we consider these mappings to be TPs in our study.